# A Circuit-QED Lattice System with Flexible Connectivity and Gapped Flat Bands for Photon-Mediated Spin Models

### v1

@@ III.2.2 CPW Arrays
where $\omega_{\mu}$ is the resonant frequency of a single resonator, $\langle n,n^{\prime}\rangle$ denotes nearest-neighbor pairs of resonators, and $t^{(\mu)}_{n,n^{\prime}}$ denotes the effective hopping strength between any two resonators participating in a 3-way coupler. Here we consider only the case where all coupling capacitors are equal and all resonators have identical resonance frequencies. (The more general case is derived in Ref. [46].) A tight-binding Hamiltonian of this form is extremely general; however, the versions realized in CPW lattices have several universal and distinctive features. First, because a coupling capacitor is equivalent to a modulation of the kinetic energy rather than a potential energy barrier, the energetics of the coupled modes are unusual, with long-wavelength normal modes occurring at high energy and short-wavelength ones at low energy [11, 46, 12]. Second, the hopping strength is highly frequency dependent [46, 12]. Third, the half-wave and full-wave modes often give rise to different band structures, and in general two separate tight-binding models, with distinct $t^{(\mu)}_{n,n^{\prime}}$, are required to describe both sets of normal modes in a CPW lattice [18, 12].
@@
where $\tilde{\Phi}_{n,\mu}$ denotes the value of the mode function $\Phi_{\mu}(x)$ in the $n^{\mathrm{th}}$ resonator evaluated at the end participating in the coupling, and $\omega_{\mu}$ is the frequency of a single resonator. For full-wave modes, the sign of $\Phi_{2}(x)$ is the same at both ends of the resonator, so the product $\tilde{\Phi}_{n,\mu}\tilde{\Phi}_{n^{\prime},\mu}$ is always positive, leading to uniform negative values of $t^{(2)}_{n,n^{\prime}}$, which we denote by $t_{2}$. For half-wave modes, on the other hand, $\Phi_{1}(x)$ has opposite sign at opposite ends of the resonator. As a result, it is not always possible to choose a definition of $\Phi_{1}(x)$ on each resonator such that $\tilde{\Phi}_{n,1}\tilde{\Phi}_{n^{\prime},1}$ is always positive. In these cases, $t^{(1)}_{n,n^{\prime}}=\pm t_{1}$ will have constant magnitude but mixed signs. Remarkably, Ref. [12] showed that for the symmetric three-way couplers utilized to make the CPW lattice presented in Section IV, the two resulting band structures formed by $H_{1}$ and $H_{2}$ both exhibit flat bands at $\omega_{\mu}-2|t_{\mu}|$, and the remaining bands are related by a band inversion.
III.2.3 Frequency-Dependent Hopping
The form of the hopping coefficient $t_{\mu}$ has frequency dependence both in $\omega_{\mu}$ and in the mode functions $\tilde{\Phi}_{n,\mu}$. The exact mode functions for a CPW with a coupling capacitor are derived in Ref. [58, 46], but to lowest order, they can be approximated by the isolated-resonator mode functions in Eq. 11 in order to obtain a simple relation between $t_{\mu}$ and the fundamental properties of the CPW:
$t_{\mu}\approx-\frac{1}{\pi}\omega_{1}\omega_{\mu}C_{c}Z_{0}.$ (14)
where $Z_{0}$ is the impedance of the CPW. (See Appendix B.2 for details.) From this form, it follows that the hopping at the full-wave modes, $|t_{2}|$, is twice as large as the hopping at the half-wave modes, $|t_{1}|$.
However, as the ratio $t_{\mu}/\omega_{1}$ increases, additional corrections begin to appear. The hopping strengths in the device presented here have $t_{1}/\omega_{1}\approx 1\%$, and are just large enough that shifts in the band edges beyond the simple tight-binding approximation start to become visible. To simulate these effects, we use a self-consistent approximation that the effective hopping rate $t_{\mu}$ is proportional to the normal-mode frequency, rather than the fixed bare frequency of a single isolated resonator.
@@ III.2.3 Frequency-Dependent Hopping
In principle, this approach could be carried out to all orders to find a true self-consistent frequency. However, for the values of $t_{\mu}$ used here, higher-order corrections in $t_{\mu}/\omega_{\mu}$ are too small to be observed in the presence of experimental disorder in CPW lattices [11]. We therefore consider only the first-order correction in Eq. 17. This approximation, taking into account the frequency dependent impedance of the couplers, more accurately describes the normal modes of a CPW resonator array than the pure tight-binding model. In Section V.2, we use this model, applied to spectroscopic measurements of the bands of the lattice, in order to extract $\omega_{\mu}$ and $t_{\mu}^{(0)}$ for the half-wave and full-wave modes of our resonators.
III.2.4 Frequency-Dependent Coupling
Conventionally, the coupling strength $g$, as introduced in Eq. 1, between a transmon and CPW resonator is written out in terms of the appropriate transmon charge matrix element and the RMS voltage in the resonator [56]. It is uncommon to see the frequency dependence of the coupling explicitly included in this definition. However, when working with higher harmonics of CPW lattices, it is crucial to understand how the coupling is affected when operating with different modes. From Ref. [46], the coupling strength for two resonators with frequencies $\omega_{1}$ and $\omega_{2}$ scales as $\sqrt{\omega_{1}\omega_{2}}$, a fact which was used to derive the hopping between two identical CPWs in Eq. 13. Intuitively, this extends to a transmon coupled to a CPW resonator such that $g$ scales as $\sqrt{\omega_{q}\omega_{\mu}}$, since a transmon is a weakly anharmonic oscillator. Alternatively, this frequency dependence can be explicitly worked out from the standard charge matrix element and zero-point voltage form used in circuit QED [56].
@@ III.2.4 Frequency-Dependent Coupling
IV Device Design
The CPW resonator lattice platform can host a wide range of potential connectivities, with the associated band structures setting the geometric dependence of the qubit-qubit interactions. Unlike previous experiments [7] which looked only at simple 1D chains and quadratic band edges which give rise to exponentially localized qubit-qubit interactions, CPW lattices natively host quadratic, linear, and flat bands, and a much larger variety of interactions. The flat bands produced in these lattices are generalized versions of the well-known Kagome-lattice flat band which consist entirely of oscillatory states which flip sign between neighboring sites [12]. The combination of the oscillatory photonic wave functions and the triangular plaquettes of the lattice can lead to a short-range frustrated magnetic interaction. Similar results have been shown theoretically for 1D sawtooth lattice [8], but CPW lattices provide a route to extending this type of behavior to 2D and non-Euclidean cases.
IV.1 Quasi 1D lattice with gapped flat bands
Our quasi-1D lattice consists of rhombus-shaped unit cells arranged in a 1D chain. Four unit cells of our target lattice are depicted in Fig. 1(a). All lattice sites have equal resonant frequencies (on-site energies), and all nearest-neighbor pairs have equal hopping strengths, up to experimental disorder [10]. By utilizing resonators with different shapes and harnessing the flexible properties of the CPW resonator, this quasi-1D pattern can be produced on a square chip while ensuring that each resonator has the same frequency and preserving the 1D translation symmetry.
The band structure that arises from the coupled half-wave modes is shown in Fig. 1(b). There are two distinct sets of bands present; the higher set features a linear band crossing with a flat band in the middle while the lower set features a dispersive band with a quadratic band edge that touches two degenerate flat bands. The model of the band structure depicted here represents a diagonalization of the tight-binding Hamiltonian in Eq. 12 without taking into account the frequency-dependent hopping correction from Section III.2.3.
As the sign of the hopping parameters depends on the parity of the mode function (Eq. 13), the full-wave modes of this lattice have their own separate band structure [12], depicted in Fig. 1(c). The highest band is an isolated dispersive band with a quadratic band edge and the middle set of bands has another linear band crossing around a flat band. Lastly, the lowest bands are two doubly-degenerate gapped flat bands. Such gapped flat bands are extremely rare in general lattices but common in CPW lattices, where they arise from non-bipartite resonator connectivity [18, 12].
IV.2 Hardware Implementation
Figure 2: Device Design. (a) Photograph of the device after the photolithography stage. Dark regions are tantalum, and white regions are the underlying $25.4\times 25.4\ \mathrm{mm}^{2}$ sapphire substrate. (b) CAD of the lattice device with color coding to illustrate different aspects of the design. The resonators forming the lattice are shown in black, with the exception of the central unit cell, which is highlighted in red. Input/output ports in the lower left and upper right of the device allow microwave transmission measurements. The locations of the three qubits are highlighted by the tan squares, and the associated on-chip flux bias lines, which allow independent frequency control of the qubits, are shown in orange. The qubits are referred to as Q1, Q2, and Q3 from right to left. Transmon qubits, with the design shown in (d), are incorporated into the lattice using pockets opened in the side of the CPW shown in purple. To mitigate disorder, the capacitor paddles of a transmon are included in all 54 resonators, but the Josephson junctions needed to complete the transmons are only included at the three highlighted sites. (c) Effective tight-binding model describing photons in the device overlaid on the resonator network. (d) CAD design of the transmon qubit. The capacitor is digitated (yellow and dark blue regions), and the e-beam lithography pattern for the Josephson junctions is shown in red. The transmon is sandwiched between the center pin of the resonator just to its left (light blue) and its associated flux-bias line, the end of which can be seen in the lower right.
The device is fabricated on a $25.4\times 25.4\ \mathrm{mm}^{2}$ sapphire wafer. The CPW resonator lattice is then patterned on by using photolithography on a $200$ nm film of superconducting tantalum. The device after photolithography is shown in Fig. 2(a) and the CAD used during the fabrication process is displayed in Fig. 2(b). The device consists of nine unit cells of the quasi-1D lattice described in Section IV.1. The starting and ending unit cells are connected to input and output ports through which the device can be probed at microwave frequencies. (See Appendix B for details.) Each unit cell consists of six CPW resonators, highlighted in red in Fig. 2(b). The resonators in the lattice are butt coupled using 3-way capacitors [11, 18]. To maintain uniform hopping, the coupling capacitors all have $120^{\circ}$ rotation symmetry, regardless of the orientation of the resonators, see Fig. 10 for a more zoomed-in view. Fig. 2(c) shows an overlay of the lattice structure from Fig. 1(a) on top of a CAD image of the device to illustrate how the coupling scheme of the target lattice is physically realized by the hardware. The lattice is laid out in an s-curve shape in order to accomodate a long 1D lattice on a square chip and uses four different resonator shapes to optimize packing density. This process of combining resonators with different form factors to produce a single lattice with equivalent sites, while in principle avoidable in this case, is intrinsically necessary for realizing fully general devices, such as non-Euclidean lattices, on chip.
Unlike previous realizations of CPW lattices [11, 10, 18], this device incorporates transmon qubits into the resonator array. The design includes on-chip flux bias lines for up to four transmons capacitively coupled to lattice sites, three of which survived the fabrication process. The locations of the three transmons are highlighted in yellow in Fig. 2(b). We label the qubits as Q1, Q2, and Q3 from right to left. Each qubit has a corresponding on-chip flux bias line routed to it, colored orange in Fig. 2(b). DC current is applied via these lines to change the magnetic flux through the SQUID loops in the qubits, tuning their frequencies. The CAD for these flux-tunable transmons is shown in Fig. 2(d). The tantalum capacitor paddles, depicted in dark blue and yellow, are defined using photolithography, along with the CPWs.
In contrast, the SQUID loop (shown in red in Fig. 2(d)) which completes the transmon is included only in the highlighted sites. The SQUID and its constituent Al/Al2O3/Al Josephson junctions are written using electron-beam lithography and deposited using double-angle evaporation with in-situ oxidization. More detailed information on the fabrication process is presented in Appendix F.
V Characterization of the Photonic Bands
In this section, we show how the modes of the resonator lattice can be located and characterized in situ using a combination of standard transmission spectroscopy and a novel implementation of two-tone spectroscopy, referred to as mode-mode spectroscopy.
V.1 Transmission Measurements
The primary method for characterizing the photonic modes of the lattice is direct measurement of the transmission of microwave signals through the device versus frequency. In this method, the energy bands appear as frequency regions that are densely populated with high-quality-factor resonances. By comparing transmission spectroscopy of the bands with the structure of the target lattice, we confirm that the lattice maintains its structure in the presence of the incorporated qubits. Additionally, when the frequency of a qubit is tuned through the bands, the resonances that shift in response differentiate the true lattice modes from parasitic modes in the packaging that share negligible coupling with the qubits. These characterization techniques are utilized for the half-wave modes (at $\sim 5$ GHz) and the full-wave modes (at $\sim 10$ GHz) to demonstrate that their band structures are consistent with Figs. 1(b) and (c) respectively.
A full transmission spectroscopy characterization of the half-wave bands is shown in Fig. 3(a). Using the on-chip flux bias lines, the frequency of Q3 is swept from above the half-wave modes to its frequency minimum at half flux and back again, while the other two qubits, Q1 and Q2, are held at a fixed frequency near $10$ GHz, far detuned from the half-wave modes. A transmission measurement is taken at each bias point to monitor how the features respond to the passage of the qubit. Both times that Q3 crosses the half-wave bands, it leaves a trail of avoided crossings with the photonic modes while extraneous packaging modes are unaffected. Comparing transmission spectroscopy of the half-wave modes with the predicted band structure from Fig. 1(b), the lower set of bands lies between $\sim 4.81$ GHz and $\sim 4.84$ GHz while the higher set of bands lies between $\sim 4.89$ GHz and $\sim 5.04$ GHz, with the two sets separated by a bandgap that is roughly $50$ MHz wide. Equivalent transmission scans for all three qubits are shown in Fig. 6 in Appendix A. All three qubits show clear avoided crossings with modes in these two regions, although the coupling strengths to each individual lattice mode vary due to the different locations of the three qubits in the lattice.
Comparing the photonic modes from the transmission spectroscopy in Fig. 4(a) to the predicted band structure from Fig. 1(c), the middle three bands (purple, navy, and gold in Fig. 1(c)), which lie between $9.6$ GHz and $9.9$ GHz, are most clearly visible. The remaining bands, the doubly-degenerate gapped flat band and the highest energy dispersive band, are difficult to locate in transmission due to their low coupling to the input and output ports of the device. With the appropriate power settings and a high frequency resolution, it is possible to view some of the modes in these bands in transmission (see Appendix A), but the bands can be much more cleanly visualized using a new non-linear measurement technique that we term mode-mode spectroscopy, and which will be described in detail in Section V.2.
The well-resolved avoided crossings between Q3 and the photonic energy bands in Fig. 3(a) and Fig. 4(a) indicate that the device has achieved strong qubit-mode coupling, despite the fact that the photonic modes are spatially extended throughout the lattice instead of confined to a single resonator. One consequence of this is that two-tone spectroscopy of Q3 is feasible in this multimode environment. Viable modes for usage in two-tone spectroscopy can be identified from these transmission scans by selecting the individual modes that display the largest response to sweeping the qubit frequency. For example, from Fig. 3(a), the mode near $4.99$ GHz exhibits a significant frequency shift in response to the proximity of Q3 and functions well as a monitor mode. Viable monitor modes for performing two-tone spectroscopy on Q1 and Q2 are determined from the transmission scans in Appendix A.
Figure 3: Half-Wave Characterization. (a) Microwave transmission through the lattice as a function of flux applied to Q3, while the other two qubits are far detuned from the bands. The qubit Q3 is tuned through the half-wave modes to its frequency minimum at half flux ($\phi_{0}=0.5)$, creating a series of avoided crossings. (b) Non-linear mode-mode spectroscopy of the half-wave modes, taken with all three qubits far detuned from the bands. Because this measurement relies on coupling to the qubits, it is insensitive to any weakly coupled parasitic modes in the device packaging. As a result, the band gap and two sets of bands are more clearly visible here than in the transmission data. (c) Expected density of states from the band structure of an infinite lattice and empirically determined values of the hopping parameter ($|t_{1}|=2\pi\times 40$ MHz) and band center ($\omega_{1}=2\pi\times 4.889$ GHz). Dispersive bands are shown in light blue, and the flat bands in dark blue. The full height of the flat-band peaks is well above the scale shown here. The two regions in which avoided crossings are visible in (a) correspond to the two frequency regions with non-zero density of states, and the mode-mode spectroscopy shown in (b) is in good agreement with the theoretical density of states.
Figure 4: Full-Wave Characterization. Combined characterization data for the full-wave modes of the device. (a) Microwave transmission through the lattice as function of flux applied to Q3, while the other two qubits are held at fixed detuning. Near integer flux, the qubit Q3 reaches its maximum frequency and tunes through the full-wave modes, creating a series of avoided crossings. In this frequency region, unwanted package modes are more significant than at the lower-frequency modes, resulting in a rolling background in transmission and causing many lattice modes to appear as Fano resonances. Also shown in (a), using a navy-gold color scale, is non-linear spectroscopy of the qubit transition itself. Two regions of this data are overlaid on top of the transmission, revealing qubit-like mid-gap states in which the qubit is heavily dressed by the the lattice modes. (b) Non-linear mode-mode spectroscopy of the full-wave modes, taken with all three qubits far detuned from the bands. The three sets of bands expected for this set of modes are easily distinguishable. Modes in the upper quadratic band and highly localized modes in the gapped flat bands that are difficult to observe in transmission are clearly visible in this mode-mode spectroscopy. (c) Expected density of states from the band structure of an infinite lattice using empirically determined values of the hopping parameter ($|t_{2}|=2\pi\times 82$ MHz) and band center ($\omega_{2}=2\pi\times 9.726$ GHz). Dispersive bands are shown in light blue, and the flat bands in dark blue. The full height of the peaks due to the flat bands is well above the scale shown here.
V.2 Mode-Mode Spectroscopy
While transmission spectroscopy provides crucial insights about the photonic modes of the lattice, the requirement that a mode couples to both the input and output ports of the device means that many lattice modes (particularly highly localized ones) are absent from the transmission spectrum. The most dramatic example of this absence are the doubly-degenerate gapped flat bands of the full-wave modes, where the modes are nearly invisible in transmission. As a result, it is difficult to compare transmission measurements with quantitative models of the target lattice and extract experimental values for the hopping and band centers. In this section, we present a novel measurement technique for probing the lattice modes, that we refer to as mode-mode spectroscopy, which is independent of the coupling between the output port and the mode, allowing for better comparison with theoretical predictions.
In essence, mode-mode spectroscopy is an implementation of two-tone spectroscopy (as discussed in Section III.1), but is used for detecting other lattice modes in the system instead of reading out the frequency of one of the qubits. The transmission of a microwave signal through one photonic mode of the lattice (referred to as the monitor mode) is recorded while the frequency of a second pump tone is swept over the remainder of the modes. The transmons coupled to the resonator array add an element of non-linearity to the normal modes of the lattice. Therefore, when the drive tone is brought into resonance with one of the other photonic modes, the frequency of the monitor mode shifts. At low drive powers, there is a weak frequency response that originates from Kerr nonlinearity [60]. At high drive powers, a drive tone near resonance with a lattice mode "ionizes" the transmons by driving them from their ground states to one of their higher energy levels through a multiphoton process, resulting in a much larger response (see Appendix D for details) [61, 62]. The mode-mode spectroscopy we present here is all taken in this transmon ionization regime (i.e., with a very strong drive tone) to maximize the number of visible modes.
The second beneficial property of mode-mode spectroscopy is that as long as a mode can be driven, i.e., if it couples to the input port, it can be measured. Thus, some modes that are invisible in transmission, especially localized modes in the flat bands, can show up clearly in mode-mode spectroscopy. This is most dramatically evident for the full-wave modes. In the transmission spectroscopy in Fig. 4(a), neither the doubly-degenerate gapped flat bands nor the band edges of the highest energy dispersive band can be accurately characterized in transmission due to the low input or output coupling of the modes. However, using mode-mode spectroscopy (Fig. 4(b)), it is plainly apparent that the flat bands are located near $9.57$ GHz and the upper dispersive band lies between $10$ and $10.1$ GHz.
Clear identification of bands and gaps from mode-mode spectroscopy makes it possible to compare the experimental characterization of the lattice modes with a theory model of the target lattice band structure to extract the mode frequencies, $\omega_{\mu}$, and the zeroth order hopping strength, $t_{\mu}^{(0)}$, for the half-wave and full-wave modes ($\mu=1,2$). For each resonator harmonic $\mu$, we compute the infinite-system normal mode frequencies $\omega_{\mu,k}^{(1)}$ numerically using a Bloch-wave ansatz and periodic boundary conditions, incorporating the approximation outlined in Eqs. 15-17 to account for the frequency-dependent hopping of a CPW array. Fig. 3(c) and Fig. 4(c) depict the density of states calculated in this manner of the half-wave and full-wave bands. The parameters $\omega_{\mu}$ and $t_{\mu}^{(0)}$ are empirically determined by matching the band edges in the numerical density of states with the mode-mode spectroscopy data. For the half-wave bands, a hopping parameter of $|t_{1}|=2\pi\times 40$ MHz and on-site energy of $\omega_{1}=2\pi\times 4.889$ GHz were extracted. The full-wave modes yield a hopping parameter of $|t_{2}|=2\pi\times 82$ MHz and on-site energy of $\omega_{2}=2\pi\times 9.726$ GHz.
VI Qubit Characterization
VI.1 Near Resonance
Having verified the photonic modes of the device, we use conventional two-tone spectroscopy, as described in Section III.1.2, to measure the transition frequencies of the transmons embedded in the lattice and characterize their hybridization with the photonic modes. As discussed in Section V.1, monitor modes with relatively strong coupling to a qubit are preselected by identifying the modes which display the largest avoided crossings in transmission.
Direct observation of both of these bound states allows verification of two key aspects of the device. First, while $g_{0}$, the coupling between a qubit and the mode of a single isolated CPW resonator, cannot be directly measured in the multimode lattice, the frequency splitting between the two bound states provides a strong constraint. An approximate value of $g_{0}=2\pi\times 165$ MHz for the full-wave modes (and $g_{0}=2\pi\times 82.5$ MHz for the half-wave modes) is determined from comparison of the bound states and numerical simulation (see Appendix C for details).
Second, while the theoretically-predicted gapped flat bands are not conclusively identifiable in transmission, the measured flux-dependence of the lower bound state cannot be explained solely by coupling to the dispersive bands of the lattice. The bound state data therefore indicates that the qubits are interacting with strongly-coupled flat-band modes in the predicted region.
VI.2 In the band gaps: photon-mediated interactions
Bound states near the band edges, such as those in Fig. 4(a), have significant photon character, so while they can give rise to strong qubit-qubit interactions, the associated losses are unfavorable. Moving deeper into the bandgap results in significantly lower losses and a qubit-qubit interaction which is dominantly mediated by virtual photons [54]. We exploit the frequency tunability of our transmons to measure qubit-qubit interactions in this regime, mediated by either set of modes in the lattice.
To observe these interactions, one transmon is held at a fixed frequency near the bands of interest while the frequency of a second qubit is swept through resonance with it, using a feed-forward method to remove flux crosstalk (see Appendix E for details). The resultant avoided crossing due to the virtual photon exchange is mapped out with two-tone spectroscopy. This procedure is carried out at three set detunings from the bottom of the half-wave bands (Fig. 5(a)-(c)) and three detunings below the full-wave bands (Fig. 5(d)-(f)), with Q1 being tuned through resonance with Q2 in each of the six scans. In each case, as the pair of qubits are brought closer to the mediating bands, the bound states grow less localized and their photonic wave functions overlap further causing the strength of the interaction to increase. This effect is observable through the size of the avoided crossings, which grow wider as the qubit pair approaches the bands.
VII Conclusion
In this work, we present the first CPW lattice device which incorporates both a lattice with non-trivial band structure and superconducting qubits. Our results show that low disorder photonic lattices, previously characterized in resonator-only devices [11, 10, 18], survive the systematic introduction of qubits, even though transmon qubits are relatively large-area devices which can significantly perturb host resonators.
Furthermore, the generalized mode-mode spectroscopy technique developed here exploits any qubits in the lattice (and their associated Josephson nonlinearity) in order to implement fast and background-insensitive diagnostics of the photonic lattice, which were previously not possible. Previous resonator-only devices, such as those of Refs. [11, 18] fundamentally lacked the ability to distinguish CPW lattice modes from any parasitic transmission channels in the device packaging. Scanning probe methods have previously been used to measure the normal-modes of a CPW lattice directly [10]. However, these methods are extremely time-consuming and require opening up the device package to leave the full top surface available for the scanning probe, which is likely highly detrimental to the lifetimes of any qubits in the lattice. By contrast, our non-linear spectroscopy method, while not a direct spatial probe, can be carried out in a fully-closed package and strongly distinguishes between active lattice modes and package leakage.
In addition, mode-mode spectroscopy gives strong signatures as long as a mode is coupled to a qubit and can be driven. Hence, modes with weak input-output coupling and even highly-localized modes are strongly visible with this technique. Therefore, it provides a potential route to exploring the localization and delocalization of modes in a flat or hyperbolic band due to the influence of a local perturbing potential [63].
In general CPW lattices, the fundamental half-wave and second-harmonic full-wave modes exhibit distinct band structures, and only the second-harmonic band can exhibit gapped flat bands [12]. Therefore, realizing the full variety of spin-model connectivities made possible by combining photon-mediated qubit-qubit interactions and the graph-like flexibility of CPW lattices requires the ability to access both sets of modes and operate both qubits and measurements in two frequency ranges an octave apart. We demonstrate qubits which can span the required range and use conventional superconducting-qubit two-tone spectroscopy carried out via the normal modes of the lattice to observe virtual-photon-mediated interactions due to both sets of bands.
In sum, the results presented here complete the toolkit developed by previous CPW lattice experiments and demonstrate the full experimental-design process and measurements needed to realize CPW lattices with superconducting qubits, both for Euclidean and non-Euclidean lattices. This device, therefore, paves the way to driven-dissipative spin models with a large variety of connectivities. The door is now open to next-generation devices with larger numbers of qubits and 2D or hyperbolic [5] spin-model connectivities.
Acknowledgements.
Appendix A Supplementary Transmission Data

### v2

@@ A.2 CPW Arrays
where $\omega_{\mu}$ is the resonant frequency of a single resonator, $\langle n,n^{\prime}\rangle$ denotes nearest-neighbor pairs of resonators, and $t^{(\mu)}_{n,n^{\prime}}$ denotes the effective hopping strength between any two resonators participating in a 3-way coupler. Here we consider only the case where all coupling capacitors are equal and all resonators have identical resonance frequencies. (The more general case is derived in Ref. [Koch:AnnPhysBerl].) A tight-binding Hamiltonian of this form is extremely general; however, the versions realized in CPW lattices have several universal and distinctive features. First, because a coupling capacitor is equivalent to a modulation of the kinetic energy rather than a potential energy barrier, the energetics of the coupled modes are unusual, with long-wavelength normal modes occurring at high energy and short-wavelength ones at low energy [Houck:earlylattice, Koch:AnnPhysBerl, Kollar:2019linegraph]. Second, the hopping strength is highly frequency dependent [Koch:AnnPhysBerl, Kollar:2019linegraph]. Third, the half-wave and full-wave modes often give rise to different band structures, and in general two separate tight-binding models, with distinct $t^{(\mu)}_{n,n^{\prime}}$, are required to describe both sets of normal modes in a CPW lattice [Kollar:2019hyperbolic, Kollar:2019linegraph].
@@
where $\tilde{\Phi}_{n,\mu}$ denotes the value of the mode function $\Phi_{\mu}(x)$ in the $n^{\mathrm{th}}$ resonator evaluated at the end participating in the coupling, and $\omega_{\mu}$ is the frequency of a single resonator. For full-wave modes, the sign of $\Phi_{2}(x)$ is the same at both ends of the resonator, so the product $\tilde{\Phi}_{n,\mu}\tilde{\Phi}_{n^{\prime},\mu}$ is always positive, leading to uniform negative values of $t^{(2)}_{n,n^{\prime}}$, which we denote by $t_{2}$. For half-wave modes, on the other hand, $\Phi_{1}(x)$ has opposite sign at opposite ends of the resonator. As a result, it is not always possible to choose a definition of $\Phi_{1}(x)$ on each resonator such that $\tilde{\Phi}_{n,1}\tilde{\Phi}_{n^{\prime},1}$ is always positive. In these cases, $t^{(1)}_{n,n^{\prime}}=\pm t_{1}$ will have constant magnitude but mixed signs.
To convert $t_{\mu}$ into a simpler form in terms of fundamental properties of the CPW and the harmonics of the resonator, we evaluate $\tilde{\Phi}_{\mu}$ using the normalization condition
$c_{\ell}\int_{0}^{L}{\Phi_{\mu}^{2}(x)dx}=1,$
However, the form above now depends on the the capacitance per unit length $c_{\ell}$, which cannot be measured directly. A more convenient form can be obtained by eliminating $c_{\ell}$ and $L$ via their relationship to the impedance $Z_{0}$ of the CPW transmission line and the fundamental frequency $\omega_{1}$ of the resonator [pozar]: $L\sqrt{\ell_{\ell}c_{\ell}}=\pi/\omega_{1}$ and $Z_{0}=\sqrt{\ell_{\ell}/c_{\ell}}$, where $\ell_{\ell}$ is the inductance per unit length of the CPW. Hence, $2/c_{\ell}L=2\omega_{1}Z_{0}/\pi$, and
$t_{\mu}\approx-\frac{1}{\pi}\omega_{1}\omega_{\mu}C_{c}Z_{0}\approx-\frac{\mu}{\pi}\omega_{1}^{2}C_{c}Z_{0}.$ (18)
Thus, the effective hopping rate scales linearly with the coupling capacitance and mode index $\mu$, but quadratically with the fundamental mode frequency.
A.3 Frequency-Dependent Hopping
As the ratio $t_{\mu}/\omega_{1}$ increases, additional corrections to the hopping rate begin to appear. The hopping strengths in the device presented here have $t_{1}/\omega_{1}\approx 1\%$, and are just large enough that shifts in the band edges beyond the simple tight-binding approximation start to become visible. To simulate these effects, we use a first-order perturbative approximation that the effective hopping rate $t_{\mu}$ is proportional to the normal-mode frequency, rather than the fixed bare frequency of a single isolated resonator.
@@ A.3 Frequency-Dependent Hopping
In principle, this approach could be carried out to all orders to find a true self-consistent frequency. However, for the values of $t_{\mu}$ used here, higher-order corrections in $t_{\mu}/\omega_{\mu}$ are too small to be observed in the presence of experimental disorder in CPW lattices [Houck:earlylattice]. We therefore consider only the first-order correction in Eq. 21. This approximation, taking into account the frequency dependent impedance of the couplers, more accurately describes the normal modes of a CPW resonator array than the pure tight-binding model. In Section V.2, we use this model, applied to spectroscopic measurements of the bands of the lattice, in order to extract $\omega_{\mu}$ and $t_{\mu}^{(0)}$ for the half-wave and full-wave modes of our resonators.
Since transmission and mode-mode spectroscopy measurements are not momentum resolved, it is not possible to directly compare the perturbative approximation to the band structure with the experimental data. Instead, we compute the perturbed density of states in the limit of a large lattice with periodic boundary conditions and compare that to the frequency ranges in which we observe the presence or absence of modes. In particular, we optimize the theoretical parameters to match the locations of the edges of the band gaps.
A.4 Frequency-Dependent Coupling
Conventionally, the coupling strength $g$, as introduced in Eq. 1, between a transmon and CPW resonator is written out in terms of the appropriate transmon charge matrix element and the RMS voltage in the resonator [Koch:transmon]. It is uncommon to see the frequency dependence of the coupling explicitly included in this definition. However, when working with higher harmonics of CPW lattices, it is crucial to understand how the coupling is affected when operating with different modes. From Ref. [Koch:AnnPhysBerl], the coupling strength for two resonators with frequencies $\omega_{1}$ and $\omega_{2}$ scales as $\sqrt{\omega_{1}\omega_{2}}$, a fact which was used to derive the hopping between two identical CPWs in Eq. 16. Intuitively, this extends to a transmon coupled to a CPW resonator such that $g$ scales as $\sqrt{\omega_{q}\omega_{\mu}}$, since a transmon is a weakly anharmonic oscillator. Alternatively, this frequency dependence can be explicitly worked out from the standard charge matrix element and zero-point voltage form used in circuit QED [Koch:transmon].
@@ A.4 Frequency-Dependent Coupling
Appendix B Supplementary Transmission Data
Figure 6: High-power half-wave transmission. (a) Transmission data of the half-wave modes as a function of flux applied to Q1 while the remaining two qubits are detuned from the range of interest. (b)-(c) Analogous data for flux applied to Q2 and Q3, respectively. At the relatively high input power used here, even very narrow weakly-coupled modes are clearly visible. However, avoided crossings in stronger modes (such as those near $4.96$ GHz) show glitches due to non-linear effects and transmon ionization. The sharp dip in midgap transmission near $4.85$ GHz matches the frequency of a non-topological edge state predicted by finite-size tight-binding calculations featuring the exact termination of our lattice.